OpenStudy (anonymous):

2=-1^2+4*-1+1 solve ill give medal

OpenStudy (blurbendy):

is that 2 = -1^2 or 2 = (-1)^2 ?

OpenStudy (blurbendy):

also, are you missing a variable somewhere?

OpenStudy (anonymous):

its 2=-1^2+4*(-1)+1

OpenStudy (primeralph):

@blurbendy Beats me too.

OpenStudy (anonymous):

the equation is y=x^2+4x+1 and the pair is (-1,2)

OpenStudy (primeralph):

You can't really solve something without a variable.